# TD-Paint: Faster Diffusion Inpainting Through Time-aware Pixel Conditioning

### v1

@@ Abstract
Diffusion models have emerged as highly effective techniques for inpainting, however, they remain constrained by slow sampling rates. While recent advances have enhanced generation quality, they have also increased sampling time, thereby limiting scalability in real-world applications. We investigate the generative sampling process of diffusion-based inpainting models and observe that these models make minimal use of the input condition during the initial sampling steps. As a result, the sampling trajectory deviates from the data manifold, requiring complex synchronization mechanisms to realign the generation process. To address this, we propose Time-aware Diffusion Paint (TD-Paint), a novel approach that adapts the diffusion process by modeling variable noise levels at the pixel level. This technique allows the model to efficiently use known pixel values from the start, guiding the generation process toward the target manifold. By embedding this information early in the diffusion process, TD-Paint significantly accelerates sampling without compromising image quality. Unlike conventional diffusion-based inpainting models, which require a dedicated architecture or an expensive generation loop, TD-Paint achieves faster sampling times without architectural modifications. Experimental results [^1] Our code is provided in supplementary materials. across three datasets show that TD-Paint outperforms state-of-the-art diffusion models while maintaining lower complexity.
@@ 1 Introduction
Figure 1: (left) Illustration of the RePaint [1] generation process. The inpainting generation process of RePaint applies a cycle of backward ($q(x_{t}|x_{t-1})$ purple lines —) and forward ($p_{\theta}(x_{t-1}|x_{t})$ green lines —) diffusion steps to the samples. Several cycles are needed to synchronize the condition to which noise is added and the current generation. It can be observed that the intermediate steps of the generation process lack consistency, changing from a man with dark hair at $\hat{x}_{0,T-1}$ to a man with blond hair at $\hat{x}_{0,t}$ to a woman at $\hat{x}_{0,t+k}$. These changes occur due to the synchronization process, where the initially predicted images $\hat{x}_{0,t}$ are not well aligned with the given condition.(right) Illustration of the TD-Paint generation process.In comparison, TD-Paint can use a clean condition from the beginning of the inpainting process, resulting in a faster and more stable process. Note how the intermediate TD-Paint steps are consistent from one to another.
@@
Diffusion models have shown significant success as generative models [2, 3], by approximating the distribution of real images through a fixed Markov chain that transforms Gaussian noise into the real image distribution. During training, a forward diffusion process gradually adds noise to an image, and the model is trained to reverse this process, learning to denoise and recover the original image distribution. During generation, the backward diffusion iteratively removes noise from an initial Gaussian noise image. The trained model predicts and removes noise at each step, gradually refining the image until a photorealistic result is achieved.
@@
Methods proposed in the literature have investigated using the standard diffusion model [1, 4] for inpainting by combining the noisy condition with the current generation at each step. This technique has its limitations. While the textures match, it creates disharmony between the conditioning and generation parts. This disharmony comes from the fact that, during the early steps of the generation process, the condition contains a lot of noise that the model cannot leverage. Therefore, the generation moves away from the intended semantics and produces unsatisfactory results. An illustration of such inpainting can be found in Figure 1(left).
@@
To address this limitation, RePaint [1] introduced a resampling mechanism that repeats the diffusion steps multiple times, enabling the model to synchronize condition and generation better. Although RePaint produces highly faithful images, this resampling significantly slows the generation process. Indeed, RePaint requires approximately 5k steps to generate a single image, increasing time complexity. An illustration of RePaint’s generative process is shown in Figure 1 (left). Other approaches introduce additional constraints at each diffusion step[4, 5]. For example, [4] integrates a correction mechanism that encourages the diffusion path to remain close to the data manifold. This is achieved by minimizing the reconstruction error of the known image region relative to the unknown region.
@@
In contrast to models that use a noisy condition[1, 4], our Time-aware Diffusion Paint (TD-Paint) approach integrates the currently generated sample with a clean condition. Our method uses semantic information from the beginning of the generation process, resulting in a more efficient and cost-effective approach. To achieve this, TD-Paint uses time conditioning derived from the standard formulation of diffusion models. Instead of using a single scalar $t$ for the entire image, TD-Paint assigns a unique $t$ value to each pixel. The known image region is assigned a smaller $t$, indicating lower noise levels. In contrast, the region to be generated is assigned a $t$ value proportional to the current step in the generation process. An illustration of TD-Paint’s generative process is provided in Figure 1(right).
@@
•  We propose a novel noise modeling paradigm for diffusion models that allows for integrating varying noise levels into the input of diffusion models. TD-Paint exploits time conditioning in diffusion models to achieve faster and higher-quality generation.
@@ 2 Related Works
Inpainting aims to fill in a missing part of the image realistically. Traditional inpainting methods try to combine techniques to propagate texture and structural information onto the missing parts [6]. Some algorithms [7] use large image datasets and assume that the possible semantic space for missing regions is limited. In recent years, deep learning models for inpainting have made impressive progress using two types of generative models, VAEs [8] and GANs [9, 10].
@@
(a) Single-Stage Inpainting: Most single-stage methods use the context encoding setting introduced by Pathak et al. [11], with an encoder-decoder setup. A reconstruction loss (L2) ensures global structure consistency, while an adversarial loss ensures the reconstruction is realistic. Global consistency is an important consideration. CNNs are limited by a receptive field that grows slowly with network layers. Many layers are needed for information to travel from one side of the image to the other. Dilated convolution [12] has been used by [13] to increase the receptive field. Partial convolution [14] uses mask information to attend only the visible regions. The pyramid context encoder [15] learns an affinity map between regions in a pyramidal fashion. Fourier convolution [16] aims to provide a global receptive field to both the inpainting network and the loss function. Fast Fourier Convolutions have an image-wide receptive field, which helps with large missing areas. Mask-Aware Transformer (MAT) [17] is a transformer-based architecture that allows the processing of high-resolution images. A customized transformer block considers only valid tokens, and a style manipulation module updates convolution weights with noise to produce diverse outputs.
@@
(b) Progressive Image Inpainting: These methods seek to address global consistency by using coarse-to-fine multi-stage approaches. Multiple generations are possible for a large missing region in single-stage training. Some may have a large pixel-to-pixel distance from the original ground truth, which can be misleading when training models with pixel-wise distance losses. To mitigate this problem, Yun et al. [18] proposed a two-stage generative approach, where the first stage optimizes a coarse result using L1 loss with spatial discount, and the second stage further refines the result using global and local critics. Gated convolution [19] has been used in a coarse-to-fine network to learn valid pixels [20].
@@
(c) Prior Knowledge Inpainting: These methods leverage and mine information from generative models. Prior Guided GAN (PGG) [16] uses the latent space of a pre-trained GAN and learns to map masked images to this space using an encoder. A masked image can be mapped to a latent code during inference, and the generator can produce a corresponding inpainted image. Deep Generative Prior [21] relaxes the frozen generator assumption of GAN inversion methods and proposes progressively refining each layer. PSP [22] uses StyleGAN [23] latent space to encode an image into its latent space. Inpainting is formulated as a domain translation task performed in the latent space of StyleGAN, removing adversarial components from the training process.
@@
(d) Diffusion Model Inpainting: While GANs have recently shown impressive results, most applications are limited to generating a specific domain. In contrast, Diffusion models have gained traction for image generation; Denoising Diffusion Probabilistic Model (DDPM) [24] and Denoising Diffusion Implicit Models (DDIM) [25] can generate very diverse and high-quality images. Some unconditional diffusion models have shown the ability to perform zero-shot inpainting [26, 27] but provide only qualitative results. The Pixel Spread Model (PSM) [5] uses a decoupled probabilistic model that combines the efficiency of GAN optimization with the prediction traceability of diffusion models. It also employs a specific design architecture. Latent Diffusion Models (LDM) [3] decouple two tasks: image processing and compression, and the denoising process is learned in latent space instead of pixel space. Inpainting is performed by encoding the masked input image, downsampling the inpainting mask, and concatenating them as additional conditions to the denoising model. Designing an image and mask conditional diffusion model requires a special architecture to accept additional input for inpainting, as done in [3, 5], and often treats inpainting as a domain translation task. In contrast, TD-Paint does not require any architecture modification by directly combining the masked input and current generation with different noise levels. The Manifold Constrained Gradient (MCG) [4] adds a correction term to ensure each sample step remains close to the data manifold, allowing for more stable inpainting. RePaint [1] is close to our work. It combines the noisy condition region with the current generation. The diffusion model updates the missing pixel based on the surrounding context. We observe that during the early inpainting step of RePaint, the condition is dominated by noise and does not contain any semantic information. This causes the model prediction to deviate from the target manifold (see Figure 1(left)). A resampling mechanism is needed to synchronize the condition and generation regions, allowing semantically corrected images to be produced at the cost of significantly increasing computation time.
@@
Instead of degrading the condition to the same level as the generation, we propose keeping it clean and conditioning the missing part on the known pixels from the beginning of the generation process (see Figure 1(right)). Although it requires fine-tuning the model, our approach provides much faster sampling by avoiding resampling [1] or additional constraint [4]. We use temporal information to model the amount of detail in the image. The condition, which contains clean data, is assigned a low noise level while the generation region starts with maximum noise that gradually decreases during the process. This approach allows for a direct combination of the clean and generated regions in the same space without changing the model architecture.
@@ 3 Background On Inpainting Diffusion Models
The reverse denoising process $p_{\theta}(x_{t-1}|x_{t},t)$ allows to generate from the data distribution by first sampling from $x_{T}\sim\mathcal{N}(0,\mathbf{I})$ and iteratively reducing the noise in the sequence $x_{T},...,x_{0}$. The model $\epsilon_{\theta}(x_{t},t)$ is trained to predict the added noise $\epsilon$ to produce the sample $x_{t}$ at time step $t$. The model is trained using mean square error (MSE):
@@ 4.1 Noise Modeling for Fast Inpainting
The final component of TD-Paint training is the strategy used for $\phi_{\text{train}}$. A random patch size and a proportion of patches are sampled to define a condition. The input image is then divided into known and unknown regions based on this sampled proportion. The possible patch sizes are defined as powers of two, i.e., ${2^{i}|i\in\mathbb{N},2^{i}\leq w}$, up to the maximum size of the image. The fraction of pixels designated as the known region is represented by a real value within the interval $[0,1]$, ensuring that at least one patch remains in the unknown region. For example, as illustrated in Figure 2 for a patch size of $128$, $25\%$ of the pixels are assigned to the known region.
@@
5:     $x^{\text{$\oplus$}}_{\phi_{t}}=\sqrt{\bar{\alpha}_{\phi_{t}}}x^{\text{$\oplus$}}+\sqrt{1-\bar{\alpha}_{\phi_{t}}}\epsilon$
6:     $x^{\text{$\ominus$}}_{\phi_{t}}=\sqrt{\bar{\alpha}_{\phi_{t}}}x_{t-1}+\sqrt{1-\bar{\alpha}_{\phi_{t}}}\epsilon$
7:     $\tau={\phi_{t}}\odot(1-m)+{\phi_{t}}\odot m$
8:     $x_{\tau}=x^{\text{$\ominus$}}_{\phi_{t}}\odot(1-m)+x^{\text{$\oplus$}}_{\phi_{t}}\odot m$
9:     $x_{t-1}=\frac{1}{\sqrt{\bar{\alpha}_{\tau}}}\left(x_{\tau}-\frac{\beta_{\tau}}{\sqrt{1-\bar{\alpha}_{\tau}}}\epsilon_{\theta}(x_{\tau},\tau)\right)+\sigma_{\tau}z$
10:  end for
11:  return  $x_{0}$
@@ 4.2 Time-aware Diffusion Architecture
Diffusion models using U-Net architectures [28] can incorporate a time map $\tau$ without requiring architectural modifications. This adaptation is achieved by applying pixel-wise time conditioning to each pixel in the feature maps.
@@
where $h^{l}\in\mathbb{R}^{c_{l}\times h_{l}\times w_{l}}$ are the current features for layer $l$, $GN$ is a group normalization layer, while $L_{\text{scale}}$ and $L_{\text{shift}}$ are linear layers that change the dimension of $\gamma$ from $d$ to $c_{l}$. We apply the pixel-wise time conditioning across each spatial dimension of $h^{l}$ by using $\tau_{i,j}$ instead of $t$ as follows:
@@
where $\Gamma\in\mathbb{R}^{d\times h\times w}$ is the embedding of time $\tau$, and $D$ is a downscaling operation function which rescales $\tau$ to $h_{l}\times w_{l}$ [^2] We use bilinear interpolation, but min-pool or other techniques could be used with similar effect..

### v2

@@ Abstract
Diffusion models have emerged as highly effective techniques for inpainting, however, they remain constrained by slow sampling rates. While recent advances have enhanced generation quality, they have also increased sampling time, thereby limiting scalability in real-world applications. We investigate the generative sampling process of diffusion-based inpainting models and observe that these models make minimal use of the input condition during the initial sampling steps. As a result, the sampling trajectory deviates from the data manifold, requiring complex synchronization mechanisms to realign the generation process. To address this, we propose Time-aware Diffusion Paint (TD-Paint), a novel approach that adapts the diffusion process by modeling variable noise levels at the pixel level. This technique allows the model to efficiently use known pixel values from the start, guiding the generation process toward the target manifold. By embedding this information early in the diffusion process, TD-Paint significantly accelerates sampling without compromising image quality. Unlike conventional diffusion-based inpainting models, which require a dedicated architecture or an expensive generation loop, TD-Paint achieves faster sampling times without architectural modifications. Experimental results across three datasets show that TD-Paint outperforms state-of-the-art diffusion models while maintaining lower complexity. Github code: https://github.com/MaugrimEP/td-paint
@@ 1 Introduction
Figure 1: Comparison of noisy-condition models (e.g., RePaint) and TD-Paint generation processes. Reverse denoising steps ($p_{\theta}(x_{t-1}|x_{t})$) are depicted in purple lines —, forward noising steps ($q(x_{t}|x_{t-1})$) are shown in green lines —. Here, $x_{t}$ represents the input to the diffusion process at step $t$, $mask$ denotes the conditional area (shown in white) and the region to be generated (shown in black), and $\hat{x}_{t}$ represents the model’s prediction at diffusion step $t$. (left) RePaint (Lugmayr et al., 2022) generation process. RePaint applies a cycle of reverse and forward diffusion steps. It can be observed that the intermediate steps of the generation process lack consistency, changing from a man with dark hair at $\hat{x}_{T-1}$ to a man with blond hair at $\hat{x}_{t}$ to a woman at $\hat{x}_{t+k}$. These changes occur due to the synchronization process, where the initially predicted images $\hat{x}_{t}$ are not well aligned with the given condition.(right) TD-Paint generation process. In comparison, TD-Paint can use a clean condition from the beginning of the inpainting process, resulting in a faster and more stable process. Note how the intermediate TD-Paint steps are consistent from one to another.
@@
Diffusion models have shown significant success as generative models (Dhariwal & Nichol, 2021; Rombach et al., 2022), by approximating the distribution of real images through a fixed Markov chain that transforms Gaussian noise into the real image distribution. During training, a forward diffusion process gradually adds noise to an image, and the model is trained to reverse this process, learning to denoise and recover the original image distribution. During generation, the backward diffusion iteratively removes noise from an initial Gaussian noise image. The trained model predicts and removes noise at each step, gradually refining the image until a photorealistic result is achieved.
@@
Methods proposed in the literature have investigated using the standard diffusion model (Lugmayr et al., 2022; Chung et al., 2022) for inpainting by combining the noisy condition with the current generation at each step. This technique has its limitations. While the textures match, it creates disharmony between the conditioning and generation parts. This disharmony comes from the fact that, during the early steps of the generation process, the condition contains a lot of noise that the model cannot leverage. Therefore, the generation moves away from the intended semantics and produces unsatisfactory results. An illustration of such inpainting can be found in Figure 1(left).
@@
To address this limitation, RePaint (Lugmayr et al., 2022) introduced a resampling mechanism that repeats the diffusion steps multiple times, enabling the model to synchronize condition and generation better. Although RePaint produces highly faithful images, this resampling significantly slows the generation process. Indeed, RePaint requires approximately 5k steps to generate a single image, increasing time complexity. An illustration of RePaint’s generative process is shown in Figure 1 (left). Other approaches introduce additional constraints at each diffusion step(Chung et al., 2022; Li et al., 2023). For example, (Chung et al., 2022) integrates a correction mechanism that encourages the diffusion path to remain close to the data manifold. This is achieved by minimizing the reconstruction error of the known image region relative to the unknown region.
@@
In contrast to models that use a noisy condition(Lugmayr et al., 2022; Chung et al., 2022), our Time-aware Diffusion Paint (TD-Paint) approach integrates the currently generated sample with a clean condition. Our method uses semantic information from the beginning of the generation process, resulting in a more efficient and cost-effective approach. To achieve this, TD-Paint uses time conditioning derived from the standard formulation of diffusion models. Instead of using a single scalar $t$ for the entire image, TD-Paint assigns a unique $t$ value to each pixel. The known image region is assigned a smaller $t$, indicating lower noise levels. In contrast, the region to be generated is assigned a $t$ value proportional to the current step in the generation process. An illustration of TD-Paint’s generative process is provided in Figure 1(right).
@@
•  We propose a novel noise modeling paradigm for diffusion models that allows for integrating varying noise levels into the input of diffusion models for the known and unknown regions. TD-Paint exploits time conditioning in diffusion models to achieve faster and higher-quality generation.
@@ 2 Related Works
Inpainting aims to fill in a missing part of the image realistically. Traditional inpainting methods try to combine techniques to propagate texture and structural information onto the missing parts (Criminisi et al., 2004). Some algorithms (Grossauer, 2004) use large image datasets and assume that the possible semantic space for missing regions is limited. In recent years, deep learning models for inpainting have made impressive progress using two types of generative models, VAEs (Kingma & Welling, 2013) and GANs (Goodfellow et al., 2020; Mirza & Osindero, 2014).
@@
(a) Single-Stage Inpainting: Most single-stage methods use the context encoding setting introduced by Pathak et al. (Pathak et al., 2016), with an encoder-decoder setup. A reconstruction loss (L2) ensures global structure consistency, while an adversarial loss ensures the reconstruction is realistic. Global consistency is an important consideration. CNNs are limited by a receptive field that grows slowly with network layers. Many layers are needed for information to travel from one side of the image to the other. Dilated convolution (Yu & Koltun, 2015) has been used by (Iizuka et al., 2017) to increase the receptive field. Partial convolution (Liu et al., 2018) uses mask information to attend only the visible regions. The pyramid context encoder (Zeng et al., 2019) learns an affinity map between regions in a pyramidal fashion. Fourier convolution (Suvorov et al., 2022) aims to provide a global receptive field to both the inpainting network and the loss function. Fast Fourier Convolutions have an image-wide receptive field, which helps with large missing areas. Mask-Aware Transformer (MAT) (Li et al., 2022) is a transformer-based architecture that allows the processing of high-resolution images. A customized transformer block considers only valid tokens, and a style manipulation module updates convolution weights with noise to produce diverse outputs.
@@
(b) Progressive Image Inpainting: These methods seek to address global consistency by using coarse-to-fine multi-stage approaches. Multiple generations are possible for a large missing region in single-stage training. Some may have a large pixel-to-pixel distance from the original ground truth, which can be misleading when training models with pixel-wise distance losses. To address this issue, Yun et al. (Yu et al., 2018) proposed a two-stage generative approach. The first stage produces a coarse output optimized with L1 loss incorporating spatial discounting, while the second stage refines the output further using both global and local critics. Gated convolution (Dauphin et al., 2017) has been used in a coarse-to-fine network to learn valid pixels (Yu et al., 2019).
@@
(c) Prior Knowledge Inpainting: These methods leverage and mine information from generative models. Prior Guided GAN (PGG) (Suvorov et al., 2022) uses the latent space of a pre-trained GAN and learns to map masked images to this space using an encoder. A masked image can be mapped to a latent code during inference, and the generator can produce a corresponding inpainted image. Deep Generative Prior (Pan et al., 2021) relaxes the frozen generator assumption of GAN inversion methods and proposes progressively refining each layer. PSP (Richardson et al., 2021) uses StyleGAN (Karras et al., 2019) latent space to encode an image into its latent space. Inpainting is formulated as a domain translation task performed in the latent space of StyleGAN, removing adversarial components from the training process.
@@
(d) Diffusion Model Inpainting: While GANs have recently shown impressive results, most applications are limited to generating a specific domain. In contrast, Diffusion models have gained traction for image generation; Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020) and Denoising Diffusion Implicit Models (DDIM) (Song et al., 2020a) can generate very diverse and high-quality images. Some unconditional diffusion models have shown the ability to perform zero-shot inpainting (Sohl-Dickstein et al., 2015; Song et al., 2020b) but provide only qualitative results. The Pixel Spread Model (PSM) (Li et al., 2023) uses a decoupled probabilistic model that combines the efficiency of GAN optimization with the prediction traceability of diffusion models. Latent Diffusion Models (LDM) (Rombach et al., 2022) decouple two tasks: image processing and compression, and the denoising process is learned in latent space instead of pixel space. Inpainting is performed by encoding the masked input image, downsampling the inpainting mask, and concatenating them as additional conditions to the denoising model. Designing an image and mask conditional diffusion model requires a special architecture to accept additional input for inpainting, as done in (Rombach et al., 2022; Li et al., 2023), and often treats inpainting as a domain translation task. In contrast, TD-Paint does not require any architecture modification by directly combining the masked input and current generation with different noise levels. Differential Diffusion (Levin & Fried, 2023) introduces a new approach to soft-inpainting, where both the generated region and the conditional input are modified to ensure coherence across the entire image. This method uses LDM along with a strength map to focus on different image regions during each diffusion step. However, Differential Diffusion operates on noisy conditional inputs. The Manifold Constrained Gradient (MCG) (Chung et al., 2022) adds a correction term to ensure each sample step remains close to the data manifold, allowing for more stable inpainting. RePaint (Lugmayr et al., 2022) which aligns with our approach, combining the noisy conditional region with the current generation, where the diffusion model iteratively updates missing pixels using the surrounding context. We observe that during the early inpainting step of RePaint, the condition is dominated by noise and does not contain any semantic information. This causes the model prediction to deviate from the target manifold (see Figure 1(left)). A resampling mechanism is needed to synchronize the condition and generation regions, allowing semantically corrected images to be produced at the cost of significantly increasing computation time.
@@
Instead of degrading the condition to the same level as the generation, we propose keeping it clean and conditioning the missing part on the known pixels from the beginning of the generation process (see Figure 1(right)). Although it requires fine-tuning the model, our approach provides much faster sampling by avoiding resampling (Lugmayr et al., 2022) or additional constraints(Chung et al., 2022). We use temporal information to model the detail in the image. The condition, containing clean data, is assigned a low noise level while the generation region starts with maximum noise that gradually decreases during the process. This approach allows for a direct combination of the clean and generated regions in the same space without changing the model architecture.
@@ 3 Background On Inpainting Diffusion Models
The reverse denoising process $p_{\theta}(x_{t-1}|x_{t},t)$ allows generating from the data distribution by first sampling from $x_{T}\sim\mathcal{N}(0,\mathbf{I})$ and iteratively reducing the noise in the sequence $x_{T},...,x_{0}$. The model $\epsilon_{\theta}(x_{t},t)$ is trained to predict the added noise $\epsilon$ to produce the sample $x_{t}$ at time step $t$. The model is trained using mean square error (MSE):
@@ 4.1 Noise Modeling for Fast Inpainting
The final component of TD-Paint training is the strategy used for $\phi_{\text{train}}$. A random patch size and a proportion of patches are sampled to define a condition. Based on this sampled proportion, the input image is then divided into known regions $x^{\text{$\oplus$}}$ and unknown regions $x^{\text{$\ominus$}}$ (with corresponding $t^{\text{$\oplus$}}=0$ and $t^{\text{$\ominus$}}=t$). The possible patch sizes are defined as powers of two, i.e., ${2^{i}|i\in\mathbb{N},2^{i}\leq w}$, up to the maximum size of the image. The fraction of pixels designated as the known region is represented by a real value within the interval $[0,1]$, ensuring that at least one patch remains in the unknown region. For example, as illustrated in Figure 2 for a patch size of $128$, $25\%$ of the pixels are assigned to the known region.
@@
5:     $x^{\text{$\oplus$}}_{\phi_{t}^{\text{$\oplus$}}}=\sqrt{\bar{\alpha}_{\phi_{t}^{\text{$\oplus$}}}}x^{\text{$\oplus$}}+\sqrt{1-\bar{\alpha}_{\phi_{t}^{\text{$\oplus$}}}}\epsilon$
6:     $\tau={\phi_{t}}^{\text{$\ominus$}}\odot(1-m)+{\phi_{t}}^{\text{$\oplus$}}\odot m$
7:     $x_{\tau}=x^{\text{$\ominus$}}_{\phi_{t}^{\text{$\ominus$}}}\odot(1-m)+x^{\text{$\oplus$}}_{\phi_{t}^{\text{$\oplus$}}}\odot m$
8:     $x_{t-1}=\frac{1}{\sqrt{\bar{\alpha}_{\tau}}}\left(x_{\tau}-\frac{\beta_{\tau}}{\sqrt{1-\bar{\alpha}_{\tau}}}\epsilon_{\theta}(x_{\tau},\tau)\right)+\sigma_{\tau}z$
9:  end for
10:  return  $x_{0}$
@@ 4.2 Time-aware Diffusion Architecture
Diffusion models using U-Net architectures (Ronneberger et al., 2015) can incorporate a time map $\tau$ without requiring architectural modifications. This adaptation is achieved by applying pixel-wise time conditioning to each pixel in the feature maps.
@@
where $h^{l}\in\mathbb{R}^{c_{l}\times h_{l}\times w_{l}}$ are the current features for layer $l$, $GN$ is a group normalization layer, while $L_{\text{scale}}$ and $L_{\text{shift}}$ are linear layers that change the dimension of $\gamma$ from $d$ to $c_{l}$. We apply the pixel-wise time conditioning across each spatial dimension $h^{l}$ by using $\tau_{i,j}$ instead of $t$ as follows:
@@
where $\Gamma\in\mathbb{R}^{d\times h\times w}$ is the embedding of time $\tau$, and $D$ is a downscaling operation function that rescales $\tau$ to $h_{l}\times w_{l}$ [^1] We use bilinear interpolation, but min-pool or other techniques could be used with a similar effect..